OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) 1 –7 2 –10 3 –13 4 –16 Write a rule for the linear function in the table. x f(x) 1 3 2 2 3 1 4 0

OpenStudy (jlastino):

For the 1st one (1,-7) and (2,-10) y+7= (-10+7)/1 (x-1) y+7=-1 (x-1) y= -x+1-7 f(x) = -x-6 2nd (1,3) (2,2) y-3 = 2-3/2-1 (x-1) y-3 = -1/1 (x-1) y= -x+4 f(x)=-x+4

OpenStudy (anonymous):

A. f(x) = x – 4 B. f(x) = 3x + 4 C. f(x) = –3x – 4 D. f(x) = x – 7 those are the answers for the 2nd 1 which 1 is it? A. f(x) = 4x B. f(x) = x – 4 C. f(x) = x + 4 D. f(x) = 4 – x and for the other 1

OpenStudy (phi):

Do you know how to find the equation of a line if you have 2 points? If you do, that is how you get the answer.

OpenStudy (anonymous):

nope i dont lol i tryed and been on it for hours i really dont

OpenStudy (jlastino):

The answer for the 2nd one is C f(x)=-3x-4 i made a mistake sorry for the 1st one its D f(x)= 4-x

OpenStudy (anonymous):

thankyou